Find the measure of an angle if its measure is 5degrees less than four times the measure of its complement
Found 2 solutions by unlockmath, ankor@dixie-net.com:
Answer by unlockmath(1688)

Hello,
We know that Complementary angles equal 90 degrees, therefore we can set up the following equation:
Let x represent the angle.
x+4x-5=90
Combine and add 5
5x=95
x=19 degrees
The other angle equals:
71 degrees.
Make sense?
RJ
www.math-unlock.com

Answer by ankor@dixie-net.com(22740)

Find the measure of an angle if its measure is 5degrees less than four times the measure of its complement
:
a + 5 = 4(90 - a)
a + 5 = 360 - 4a
a + 4a = 360 - 5
5a = 355
a = 355/5
a = 71 degrees